# Transformer Block Coupling and its Correlation with Generalization in LLMs

### v2

@@ Abstract
Large Language Models (LLMs) have made significant strides in natural language processing, and a precise understanding of the internal mechanisms driving their success is essential. In this work, we trace the trajectories of individual tokens as they pass through transformer blocks, and linearize the system along these trajectories through their Jacobian matrices. By examining the relationships between these Jacobians, we uncover a transformer block coupling phenomenon in a variety of LLMs, characterized by the coupling of their top singular vectors across tokens and depth. Our findings reveal that coupling positively correlates with model performance, and that this relationship is stronger than with other hyperparameters, namely parameter budget, model depth, and embedding dimension. We further investigate the emergence of these properties through training, noting the development of coupling, as well as an increase in linearity and layer-wise exponential growth in the token trajectories. These collective insights provide a novel perspective on the interactions between token embeddings, and prompt further approaches to study training and generalization in LLMs.
@@ 1 Introduction
In recent years, many openly available Large Language Models (LLMs) have been released, achieving state-of-the-art results on task-specific benchmarks. The abundance of models, each with differing architecture and training methodology, motivates comparing the underlying mechanisms that drive generalization.
Transformers (Vaswani et al., 2017) can be represented as discrete, nonlinear, coupled dynamical systems, operating in high dimensions (Greff et al., 2016; Papyan et al., 2017; Haber & Ruthotto, 2017; Ee, 2017; Ebski et al., 2018; Chen et al., 2018; Bai et al., 2019; Rothauge et al., 2019; Gai & Zhang, 2021; Li & Papyan, 2023). The term discrete reflects the network’s finite depth; nonlinear refers to the model’s nonlinear components; coupled refers to the interdependent token trajectories enabled by self-attention, and dynamical is due to the residual connections spanning various layers.
Throughout this work, we investigate whether we can identify structural characteristics that consistently occur in pretrained LLMs that are related to their generalization, as measured on the HuggingFace OpenLLM Leaderboard benchmark. To this end, we study and quantify the trajectories of token embeddings and their interactions, through linearizations of LLM layers given by the Jacobians of the associated transformer blocks.
This perspective raises several questions:
Q1. What regularity properties do these trajectories exhibit individually, and what are their relations with one another? More concretely, what is the relation between the Jacobians across different tokens and transformer layers?
@@
Figure 1: Relationship between Coupling and Benchmark Score. The plot illustrates the relationship between average coupling (negative mis-coupling) and benchmark score across various base LLMs (that are not fine-tuned). The data shows a noticeable trend: increased alignment correlates with improved benchmark scores. The regression fit produces an $R^{2}$ value of $0.7214$.
@@ 1.1 Contributions
1. Coupling. The singular vectors of the Jacobians of transformer blocks couple across depth (Figures LABEL:fig:paris_all_8to16, LABEL:fig:paris_all_8to16_appendix) and tokens (Figures LABEL:fig:coupling_inoutsame, LABEL:fig:coupling_fixedin, LABEL:fig:coupling_fixedout, LABEL:fig:coupling_inoutsame_appendix, LABEL:fig:coupling_fixed_input_appendix, LABEL:fig:coupling_fixed_output_appendix). Further, coupling across Jacobians emerges with training (Figures 2, LABEL:fig:paris_all_8to16, LABEL:fig:coupling_inoutsame, LABEL:fig:coupling_fixedin, LABEL:fig:coupling_fixedout).
2. Generalization. The strength of coupling is correlated with benchmark performance on the HuggingFace Open LLM Leaderboard (Beeching et al., 2023) (Figures 1, LABEL:fig:alignment_vary_k). Additionally, coupling is more strongly correlated with generalization than parameter budget, model depth, and token embedding dimension (Figure LABEL:fig:correlation_hyperparams).
3. Regularity. Linearity in hidden trajectories emerges with training (Figures LABEL:figs:lss-training, LABEL:figs:pca, 13, 15, 23, 24). Exponential growth occurs in contiguous token representations as a function of depth (Figures 14, 22, 23, LABEL:figs:expo-training).
We provide a new perspective on token embedding interactions within LLMs by examining transformers layers through their Jacobians. Our results display the effect of training on transformer blocks, and suggest potential approaches for promoting generalization in LLMs.
Figure 2: Emergence of Singular Vector Coupling with Training. Normalized coupling (Section 3.1) evaluated for $J_{i,j}^{d_{1}},J_{i,j}^{d_{2}}$, with $|d_{1}-d_{2}|\leq 5$. Measurements are taken for checkpoints of Pythia 12B, Pythia 6.9B, Pythia 1B (Biderman et al., 2023) at training steps $\{1,16,128,13\text{k},23\text{k},33\text{k},\ldots,143\text{k}\}$. Coupling is low at early training iterations, and increases with training, with Pythia 12B and 6.9B displaying greater coupling than Pythia 1B in final training steps.
@@ 2 Background on Large Language Models
We describe LLMs as a deep composition of functions that iteratively transform token embeddings. In the input layer, $l=0$, textual prompts undergo tokenization and are combined with positional encodings to create an initial high-dimensional embedding, denoted by $x_{i}^{0}\in\mathbb{R}^{d_{\text{model}}}$ for the $i^{\text{th}}$ token. When these embeddings are stacked, they form a matrix:
@@
The output $X^{L}$ from the final block $F^{L}$ is passed into a bias-free linear layer $M\in\mathbb{R}^{d_{\text{vocab}}\times d_{\text{model}}}$, with $d_{\text{vocab}}$ denoting the size of the token vocabulary and $d_{\text{model}}$ is the dimension of the token embeddings. This layer $M$ computes final-layer logits for each token embedding, $\ell_{i}=Mx_{i}^{L}$. The prediction for the next token is then determined by selecting the maximal logit value: $\arg\max_{v\in\text{tokens}}\ell_{v,n}.$
@@ 3.1 Coupling of Singular Vectors of Jacobians
Jacobians. Coupling is investigated through analyzing the linearizations of transformer blocks given by their Jacobians
@@
defined for each layer $l\in\{1,\ldots,L\}$, and pair of tokens $t_{1},t_{2}\in\{1,\ldots,n\}$. Note that this is the Jacobian matrix for each transformer block without the contribution from the skip connection from the input of the block, similar to the quantity measured by Li & Papyan (2023), but extended for cross-token interactions. Also, due to the causal structure of the representations, $J_{t_{1}t_{2}}^{l}=0$ if $t_{1}>t_{2}$, so we restrict our attention to the case $t_{1}\leq t_{2}$.
Singular value decomposition. The SVD of the Jacobians $J_{t_{1}t_{2}}^{l}$ are computed, i.e., $J_{t_{1}t_{2}}^{l}=U_{l}S_{l}V_{l}^{\top}$ (with superscripts $t_{1},t_{2}$, indicating the tokens, omitted for clarity), where $U_{l}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{model}}}$ and $V_{l}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{model}}}$ are matrices of left and right singular vectors respectively, and $S_{l}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{model}}}$ is the singular value matrix.
To measure coupling between Jacobians $J_{t_{1}t_{2}}^{l}$ and $J_{t_{1}^{\prime}t_{2}^{\prime}}^{l^{\prime}}$, define $A_{t_{1}t_{2}t_{1}^{\prime}t_{2}^{\prime}}^{ll^{\prime},K}\coloneq U_{l^{\prime},K}^{\top}J_{t_{1}t_{2}}^{l}V_{l^{\prime},K}$ for $l,l^{\prime}\in\{1,\ldots,L\}$ and $t_{1},t_{2},t_{1}^{\prime},t_{2}^{\prime}\in\{1,\ldots,n\}$, where $U_{l^{\prime},K}$ and $V_{l^{\prime},K}$ are the sub-matrices with columns that are the top-$K$ left and right singular vectors of $J_{t_{1}^{\prime}t_{2}^{\prime}}^{l^{\prime}}$, respectively.
Figure 3: Transformer Block Coupling. Displayed is a visualization of the various types of transformer block coupling, along with brief instructions on computing both the Jacobians $J$ and coupling matrices $A$.
Depth-wise Coupling. To analyze coupling across transformer blocks, we fix $t$ (usually as the last token), and check alignment between $J_{tt}^{l}$ and $J_{tt}^{l^{\prime}}$ through $A_{t}^{ll^{\prime},K}$ (we write the single subscript $t$ for clarity) across layers $l,l^{\prime}\in\{1,\ldots,L\}$.
Token-wise Coupling. To study the coupling across tokens, we start by fixing two layers $l,l^{\prime}\in\{1,\ldots,L\}$, and analyze coupling across different cases of the $t_{1},t_{2},t_{1}^{\prime},t_{2}^{\prime}$. We start with the case where the input and output tokens are the same and compare $J_{tt}^{l}$ and $J_{t^{\prime}t^{\prime}}^{l^{\prime}}$ across $t,t^{\prime}\in\{1,\ldots,n\}$, which is the coupling across tokens for a token’s effect on its own trajectory. We also consider the context tokens’ impact on a trajectory by including coupling across cases with $t_{1}<t_{2}$. More specifically, we measure coupling between $J_{t_{1}t_{2}}^{l}$ and $J_{t_{1}t_{2}^{\prime}}^{l^{\prime}}$ across $t_{2},t_{2}^{\prime}\geq t_{1}$ (fixing the input token to be the same) and also between $J_{t_{1}t_{2}}^{l}$ and $J_{t_{1}^{\prime}t_{2}}^{l^{\prime}}$ across $t_{1},t_{1}^{\prime}\leq t_{2}$ (fixing the output token to be the same).
Coupling measurement. To quantify the coupling of singular vectors of the Jacobians, we measure how far the $A_{t_{1}t_{2}t_{1}^{\prime}t_{2}^{\prime}}^{ll^{\prime},K}$ are from being diagonal. More precisely, we quantify the mis-coupling of $A_{t_{1}t_{2}t_{1}^{\prime}t_{2}^{\prime}}^{ll^{\prime},K}$ using $m:\mathbb{R}^{K\times K}\to\mathbb{R}$ defined by
@@
where $\text{Diag}(A)$ is the matrix $A$ with all non-diagonal entries replaced by zero and $\lVert\cdot\rVert_{F}$ denotes the Frobenius norm. For normalized comparison between models across training in Figure 2, we normalize $A$ by their norms
$\tilde{m}(A)=\frac{\lVert A-\text{Diag}(A)\rVert_{F}}{\lVert A\rVert_{F}}$ (10)
@@ 3.3 Layer-Wise Exponential Growth
We measure the presence of exponential spacing (expodistance) of the hidden trajectories. Assuming exponential growth of the embedding norms as they flow through the hidden layers, we can estimate,
$\|x_{i}^{l}\|$ $\approx e^{\alpha l}\|x_{i}^{0}\|=e^{\alpha}\|x_{i}^{l-1}\|$ (13)
for some fixed $\alpha\in\mathbb{R}$ over all layers $l=1,\ldots L$. We quantify the validity of the representation in Equation 13 by measuring the coefficient of variation of $\alpha_{i}^{l}$, given by
@@
This measurement is motivated by the parametrization discussed in Appendix A.4, as well as empirical evidence in Figure LABEL:figs:norm-expo-traininga, and serves as a method to test the validity of the linearization presented in Equation 17.
@@ 5.1 Coupling of Jacobians Across Depth
In trained LLMs, we observe coupling of the top singular vectors of the Jacobians across depth (Figure LABEL:fig:paris_all_8to16 bottom row), evident in the low non-diagonal values with a visible diagonal present in the matrix subplots. This is consistently observed across various LLMs considered. On the other hand, in untrained models (Figure LABEL:fig:paris_all_8to16 top row), there is no coupling of Jacobians across different depths. There is coupling along the diagonal, however, because each Jacobian is trivially diagonalized by its own singular vectors. This, in addition to Figure 2, suggests that coupling across depth emerges through training.
@@ 5.2 Coupling of Jacobians Across Tokens
We also analyze the coupling of singular vectors of the Jacobians across tokens. First, we observe coupling between Jacobians where the input and output tokens are the same, that is, between $J_{l}^{tt}$ and $J_{l^{\prime}}^{t^{\prime}t^{\prime}}$ across $t,t^{\prime}\in\{1,\ldots,n\}$ (Figure LABEL:fig:coupling_inoutsame). This suggests that a token’s interaction on its own trajectory is coupled with that of other tokens. We also consider the context tokens’ interaction on a trajectory when fixing the input token $t_{1}$ (Figure LABEL:fig:coupling_fixedin), through $J_{l}^{t_{1}t_{2}}$ and $J_{l^{\prime}}^{t_{1}t_{2}^{\prime}}$ across $t_{2},t_{2}^{\prime}\geq t_{1}$, and when fixing the output token $t_{2}$ (Figure LABEL:fig:coupling_fixedout), with $J_{l}^{t_{1}t_{2}}$ and $J_{l^{\prime}}^{t_{1}^{\prime}t_{2}}$ across $t_{1},t_{1}^{\prime}\leq t_{2}$. Although generally present, the coupling in these two cases are less consistent, with variation in the strength of coupling between different token pairs. Again in untrained models (Figure LABEL:fig:coupling_inoutsame top row), there is a lack coupling of Jacobians across tokens.

### v3

@@ Abstract
Large Language Models (LLMs) have made significant strides in natural language processing, and a precise understanding of the internal mechanisms driving their success is essential. In this work, we trace the trajectories of individual tokens as they pass through transformer blocks, and linearize the system along these trajectories through their Jacobian matrices. By examining the relationships between these Jacobians, we uncover the transformer block coupling phenomenon in a multitude of LLMs, characterized by the coupling of their top singular vectors across tokens and depth. Our findings reveal that coupling positively correlates with model performance, and that this relationship is stronger than with other hyperparameters, namely parameter budget, model depth, and embedding dimension. We further investigate the emergence of these properties through training, noting the development of coupling, as well as an increase in linearity and layer-wise exponential growth in the token trajectories. These collective insights provide a novel perspective on the interactions between token embeddings, and prompt further approaches to study training and generalization in LLMs.
@@ 1 Introduction
Transformers (Vaswani et al., 2017) can be represented as discrete, nonlinear, coupled dynamical systems, operating in high dimensions (Greff et al., 2016; Papyan et al., 2017; Haber & Ruthotto, 2017; Ee, 2017; Ebski et al., 2018; Chen et al., 2018; Bai et al., 2019; Rothauge et al., 2019; Gai & Zhang, 2021; Li & Papyan, 2023). Viewing the skip connections as enabling a discrete time step, we represent the hidden representations as dynamically evolving through the layers of the network. The term nonlinear refers to the nonlinear transformations introduced by activation functions, and coupled refers to the interdependent token trajectories that interact through the MLP and self-attention blocks.
In our work, we investigate whether there are identifiable structural characteristics across 38+ pretrained LLMs, measure their emergence with training, and analyze their relationship with generalization performance. During inference, as token embeddings pass through the network, we linearize the effect of transformer blocks on the token embeddings throughout the depth of the LLM. To this end, we compute the Jacobians of distinct connections between layers or tokens, derive their singular value decompositions (SVDs), and compare the resulting singular vectors. This approach measures the degree of coupling between singular vectors to capture the operational similarity of blocks as they act on tokens. This perspective raises several questions:
Q1. What regularity properties do these trajectories exhibit, and what are their relations with one another? More concretely, what is the relation between the Jacobians across different tokens and transformer layers?
@@
Figure 1: Transformer Coupling Measurements. (a) The plot illustrates the correlation between average coupling (negative mis-coupling) and benchmark scores across base LLMs (not fine-tuned), showing that higher alignment corresponds to improved performance, with a regression fit yielding an $R^{2}$ value of 0.75 with a significant p-value of $1.56\times 10^{-6}$. (b) The mean normalized coupling (Section 3.1) is plotted as a function of training checkpoints for Pythia 12B and 6.9B (Biderman et al., 2023), measured at steps ${128,256,512,1k,2k,\ldots,128k,143k}$. (c-e) Adjacency plots illustrate the mean coupling scores between pairs of layers. Each node represents a layer, and edge weight and opacity indicate the strength of depth-wise normalized coupling. Visualizations are provided for checkpoints 1, 4k, and 143k of Pythia 12B.
@@ 1.1 Contributions
1. Coupling. The singular vectors of the Jacobians of the transformer blocks are coupled across depth (Figures 3, LABEL:fig:paris_all_8to16_appendix) and tokens (Figures 4, LABEL:fig:coupling_fixedin, LABEL:fig:coupling_fixedout, LABEL:fig:coupling_inoutsame_appendix, LABEL:fig:coupling_fixed_input_appendix, LABEL:fig:coupling_fixed_output_appendix) in several open source LLMs (Table (1)). Furthermore, coupling across Jacobians emerges with training (Figures 1b, 3, 4, LABEL:fig:coupling_fixedin, LABEL:fig:coupling_fixedout), and the coupling strength becomes more pronounced between adjacent layers with training, indicating a layer-wise locality in the interactions (Figures 1(c-e)).
2. Generalization. The strength of coupling is correlated with benchmark performance on the HuggingFace Open LLM Leaderboard (Beeching et al., 2023) (Figures 1a, LABEL:fig:alignment_vary_k). Additionally, coupling is more strongly correlated with generalization than parameter budget, model depth, and token embedding dimension (Figure LABEL:fig:correlation_hyperparams).
3. Regularity. Linearity in hidden trajectories emerges with training (Figures LABEL:figs:lss-training, LABEL:figs:pca, LABEL:fig:lss-all, 17, 24, 25), aligning with behaviour previously observed in ResNets Li & Papyan (2023). Exponential growth occurs in contiguous token representations as a function of depth (Figures LABEL:fig:expo-all, 23, 24, LABEL:figs:expo-training), starkly contrasting linear growth previously observed.
We provide a new perspective on token embedding interactions within LLMs by examining layers of a transformer through their Jacobian matrices. Our results display the effect of training on transformer blocks, and suggest potential approaches for promoting generalization in LLMs.
@@ 2 Background on Large Language Models
We describe LLMs as a deep composition of functions that iteratively transform token embeddings. In the input layer, $l=0$, the text prompts are tokenized and combined with the positional encodings to create an initial high-dimensional embedding, denoted by $x_{i}^{0}\in\mathbb{R}^{d_{\text{model}}}$ for the $i^{\text{th}}$ token. When these embeddings are stacked, they form a matrix:
@@
The output $X^{L}$ of the final block $F^{L}$ is passed into a bias-free linear layer $M\in\mathbb{R}^{d_{\text{vocab}}\times d_{\text{model}}}$, with $d_{\text{vocab}}$ denoting the size of the token vocabulary and $d_{\text{model}}$ is the dimension of the token embeddings. This layer $M$ computes final-layer logits for each token embedding, $\ell_{i}=Mx_{i}^{L}$. The prediction for the next token is then determined by selecting the maximal logit value: $\arg\max_{v\in\text{tokens}}\ell_{v,n}.$
Figure 2: Transformer Block Coupling. A visualization of the various types of transformer block coupling with brief instructions on computing both the Jacobians $J$ and coupling matrices $A$ (Section 3.1). The coupling measurement quantifies the alignment and agreement between the interactions of embeddings connections within the network. The colored subscripts in the sample matrices $A$ indicate the specific connections being compared.
@@ 3.1 Coupling of Singular Vectors of Jacobians
Jacobians. Coupling is investigated through analyzing the linearizations of transformer blocks which is given by their Jacobian matrices
@@
defined for each layer $l\in\{1,\ldots,L\}$, and pair of tokens $t_{1},t_{2}\in\{1,\ldots,n\}$. Note that this is the Jacobian matrix for each transformer block without the contribution of the skip connection from the input of the block, similar to the quantity measured by Li & Papyan (2023) which strictly analyzes the case where $t_{1}=t_{2}$.
Due to the causal structure of the representations, $J_{t_{1}t_{2}}^{l}=0$ whenever $t_{1}>t_{2}$. Hence, we restrict our attention to the case where $t_{1}\leq t_{2}$.
Singular value decomposition (SVD). We compute the SVD of the Jacobians:
$J_{t_{1}t_{2}}^{l}=U_{l}S_{l}V_{l}^{\top}$
where $U_{l},V_{l}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{model}}}$ are the matrices of left and right singular vectors respectively, and $S_{l}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{model}}}$ contains the singular values. [^1] Note that the superscripts $t_{1},t_{2}$, indicating the tokens, are omitted for clarity in the expression for the SVD.
Coupling Measurement. To measure coupling between two Jacobians
$J_{t_{1}t_{2}}^{l}=U_{l}S_{l}V_{l}^{\top}$ and $J_{t_{1}^{\prime}t_{2}^{\prime}}^{l^{\prime}}=U_{l^{\prime}}S_{l^{\prime}}V_{l^{\prime}}^{\top},$ (9)
we define the coupling matrix
$A_{t_{1}t_{2}t_{1}^{\prime}t_{2}^{\prime}}^{ll^{\prime}}$ :-$\coloneq U_{l^{\prime}}^{\top}J_{t_{1}t_{2}}^{l}V_{l^{\prime}}:$ (10)
$=U_{l^{\prime}}^{\top}U_{l}S_{l}V_{l}^{\top}V_{l^{\prime}}$ (11)
for $l,l^{\prime}\in\{1,\ldots,L\}$ and $t_{1},t_{2},t_{1}^{\prime},t_{2}^{\prime}\in\{1,\ldots,n\}$. If the singular vectors of distinct Jacobians are strongly aligned, then
$U_{l^{\prime}}^{\top}U_{l}\approx I\approx V_{l}^{\top}V_{l^{\prime}},$ (12)
implying that the coupling matrix $A$ should be strongly diagonal. Explicitly, we quantify the mis-coupling of $A$ using
@@
where $\text{Diag}(A)$ is the matrix $A$ with all non-diagonal entries replaced by zero and $\lVert\cdot\rVert_{F}$ denotes the Frobenius norm. For normalized comparison between models, we normalize by $A$;
$\tilde{m}(A)=\frac{\lVert A-\text{Diag}(A)\rVert_{F}}{\lVert A\rVert_{F}}.$ (14)
Depth-wise Coupling. To analyze coupling across transformer blocks, we fix $t$, and measure alignment between $J_{tt}^{l}$ and $J_{tt}^{l^{\prime}}$ through the matrix $A_{t}^{ll^{\prime}}$ across layers $l,l^{\prime}\in\{1,\ldots,L\}$  [^2] In the matrix $A$, we write the single subscript $t$ for clarity.
Token-wise Coupling We also quantify the coupling across tokens in several ways:
  • Self-coupling. By fixing two layers $l,l^{\prime}\in\{1,\ldots,L\}$, we analyze the case where the input and output tokens are the same. Explicity, we compare $J_{tt}^{l}$ and $J_{t^{\prime}t^{\prime}}^{l^{\prime}}$ across $t,t^{\prime}\in\{1,\ldots,n\}$, which represents the coupling across tokens for a token’s effect on its own trajectory.
  • Context Coupling. We consider the context tokens’ impact on a trajectory by measuring coupling between $J_{t_{1}t_{2}}^{l}$ and $J_{t_{1}t_{2}^{\prime}}^{l^{\prime}}$ across $t_{2},t_{2}^{\prime}\geq t_{1}$ (fixing the input token to be the same) and also between $J_{t_{1}t_{2}}^{l}$ and $J_{t_{1}^{\prime}t_{2}}^{l^{\prime}}$ across $t_{1},t_{1}^{\prime}\leq t_{2}$ (fixing the output token to be the same).
@@ 3.3 Layer-Wise Exponential Growth
We measure the presence of exponential spacing (expodistance) of the hidden trajectories. Assuming exponential growth of the embedding norms as they flow through the hidden layers, we estimate $\|x_{i}^{l}\|\approx e^{\alpha l}\|x_{i}^{0}\|=e^{\alpha}\|x_{i}^{l-1}\|$ for some fixed $\alpha\in\mathbb{R}$ over all layers $l=1,\ldots L$. We quantify the validity of this representation by measuring the coefficient of variation of $\alpha_{i}^{l}$, given by
@@
This measurement is motivated by the discussion in Section 6.1 the parametrization discussed in Appendix A.5, as well as empirical evidence in Figure LABEL:figs:norm-expo-traininga, and serves as a method to test the validity of the linearization presented in Equation 19.
3.4 Vision Transformer Training
For further investigation of coupling in transformers, we train a series of Vision Transformers (ViTs) following DEiT training (Touvron et al., 2021). We train 64 ViTs on CIFAR10 (Krizhevsky, 2009) with varied weight decay and stochastic depth rate for a fixed architecture of embedding size 192, depth 12, and 3 attention heads. Please see Appendix A.7 for further details.
Figure 3: Transformer Block Coupling across Depth. The figure shows Jacobian coupling across transformer blocks 9 to 16, using the prompt ”What is the capital of France? The capital is” to trace the final token’s trajectory. In trained models (bottom row), the diagonal pattern with minimal off-diagonal values indicates alignment of Jacobians, where top singular vectors of $J^{l^{\prime}}$ diagonalize $J^{l}$. Untrained models (top row) lack this alignment. Further details are in the Appendix. A.8 (Figure LABEL:fig:paris_all_8to16_appendix). Best viewed in color.
@@ 5.1 Coupling of Jacobians Across Depth
In trained LLMs, we observe coupling of the top singular vectors of the Jacobians across depth (Figure 3 bottom row), evident in the low non-diagonal values with a visible diagonal present in the matrix subplots. This is consistently observed across various LLMs considered. On the other hand, in untrained models (Figure 3 top row), there is no coupling of Jacobians across different depths. There is coupling along the diagonal, however, because each Jacobian is trivially diagonalized by its own singular vectors. This, in addition to Figure 1, suggests that coupling across depth emerges through training.
@@ 5.2 Coupling of Jacobians Across Tokens
We analyze the coupling of singular vectors of Jacobians across tokens. For input and output tokens that are the same ($J_{l}^{tt}$ and $J_{l^{\prime}}^{t^{\prime}t^{\prime}}$, Figure 4), we observe strong coupling, indicating that a token’s interactions along its trajectory are coupled with others. For context tokens, coupling is examined by fixing the input token ($J_{l}^{t_{1}t_{2}}$ and $J_{l^{\prime}}^{t_{1}t_{2}^{\prime}}$, Figure LABEL:fig:coupling_fixedin) or the output token ($J_{l}^{t_{1}t_{2}}$ and $J_{l^{\prime}}^{t_{1}^{\prime}t_{2}}$, Figure LABEL:fig:coupling_fixedout). While coupling exists, it is less consistent across pairs. Untrained models (Figure 4 top row) show no such coupling.
Figure 4: Transformer Block Coupling across Tokens (Same input and output tokens). The figure shows Jacobian coupling for the same input and output token across tokens, visualized using the absolute values of $A_{ll^{\prime}}^{ttt^{\prime}t^{\prime}}$ (with fixed layers $l,l^{\prime}$). In trained models (bottom row), the strong diagonal and small off-diagonal values indicate coupling, while no such coupling is present at initialization (top row). Additional details are in Appendix A.8 (Figure LABEL:fig:coupling_inoutsame_appendix).